# Perfect pulsed inline twin-beam squeezers

### v1

@@ I Introduction
Squeezed states of light are ubiquitous in quantum information and quantum computing in the optical domain Braunstein and van Loock (2005); Weedbrook et al. (2012); Braunstein (2005). Recently, there have been great advancements made on how to generate and use these states as resources in quantum computing both theoretically Christ et al. (2013, 2014); Helt and Quesada (2020); Quesada et al. (2022, 2020); Vernon et al. (2019); Arzani, Fabre, and Treps (2018); Houde and Quesada (2023) and experimentally Zhong et al. (2020, 2021, 2019); Eckstein et al. (2011); Wang et al. (2019); Deshpande et al. (2022); Arrazola et al. (2021); Triginer et al. (2020); Harder et al. (2013, 2016); Yanagimoto et al. (2023); Nehra et al. (2022).
@@
One can gain information on both sets of modes by applying a Bloch-Messiah decomposition to the Heisenberg-picture propagator describing the evolution of the squeezer Cariolaro and Pierobon (2016a, b); McCutcheon (2018); Horoshko et al. (2019). This type of decomposition is readily implemented numerically, however, it is generally very hard to obtain analytically, especially when a large number of modes are used to approximate the quasi-continuum encountered in waveguide geometries. In this paper, we look at different waveguided twin-beam squeezers, which are experimentally relevant Zhong et al. (2020, 2021, 2019), and develop a method for obtaining analytical relations between the input and output modes, without computing them explicitly. By working with discretized equations of motion and using the symmetries of the matrices describing the evolution, we are able to obtain analytical relations between the input and output modes from the Bloch-Messiah decomposition of the Heisenberg-picture propagator.
Figure 1: Schematic representation of a perfect inline squeezer. (a) Temporal mode structure with respect to central frequency $\bar{\omega}$, $f_{0}(\omega)$, that is matched to the perfect inline squeezer. (b) Wigner distribution of the injected quantum state, taken to be a single Fock state, for the matched mode. (c) Temporal mode structure with respect to central frequency $\bar{\omega}$, $f_{n}(\omega)$, of an unmatched mode. (d) Wigner distribution of the injected quantum state, taken to be a Cat state, for the unmatched mode. (e) Temporal mode structure of the output state for the matched input state. Note that the structure remains the same. (f) Wigner distribution of the output state for the matched case. The output is now is a Squeezed Fock state. (g) Temporal mode structure of the output state for the unmatched input state. Note that the structure remains the same. (h) Wigner distribution of the output state for the matched case. The Wigner distribution remains unchanged. Only the matched input generates a squeezed output.
@@
Figure 2: Waveguide poling configurations ($g(z)$) considered for the Bloch-Messiah decompositions of type-II SPDC. a) Unpoled single pass structure of length $z$, where the pump pulse enters from the left and the squeezed light exits the right. b) Poled single pass structure where the nonlinear region is domain engineered(represented by up/down arrows). The domains can be chosen such that $z_{1}=z_{2}=\cdots=z_{n}$ to generate quasi-phase-matching or more arbitrary to generate more exotic phase-matching functions (e.g. Gaussian). c) Poled double pass structure where after the first pass, the beams are reflected back into the sample by reflecting off a mirror. The polarizations of the signal and idler are also swapped by passing through a quarter-wave plate (QWP) twice. Similar to the poled single pass, the values for the $z_{i}$’s can be chosen to generate quasi-phase-matching or more exotic phase-matching functions. In all configurations we assume no reflections at any of the interfaces and no propagation losses.
In this paper we focus on non-degenerate waveguided squeezers, however, one could also consider other architectures such as resonators and degenerate squeezers. In Table 1 we summarize known results on the spectral purity and mode-matching of degenerate ring-like and waveguided structures. It has been shown that ring-like structures can be spectrally pure, however, the input and output modes are not matchedQuesada et al. (2022), just as ring-like structures that have no nonlinear response. Therefore, ring-like structures are not good candidates for perfect inline squeezers. As for degenerate waveguided squeezers, it has been shown that they are highly spectrally impureHelt and Quesada (2020). It remains to be seen if such squeezers can be mode-matched, however, simply based on the spectral purity they are not good candidates for perfect inline squeezers. In Table 2 we summarize results related to this criteria but for non-degenerate ring-like and waveguided structures. The same results and conclusions hold for non-degenerate ring-like structures. As mentioned previously, non-degenerate waveguided sources can be spectrally pure and, as we show in this paper, can also be mode-matched therefore giving rise to a perfect inline squeezer. Finally, we note that if beamsplitter operations are easy to access between the two beams of the twin-beam (as is the case for frequency degenerate type-II, for which they correspond to polarizing beamsplitters) then it is easy to show that a perfect inline squeezer can be turned into two single-mode squeezing operations acting in parallel on the two modes Kalajdzievski and Quesada (2021).
@@
The paper is organized as follows. In Sec. II we provide a detailed analysis of why it is important to match input and output modes when squeezing states other than the vacuum. In Sec. III, we give a brief description of the Bloch-Messiah decomposition as well as how to identify the input and output Schmidt modes from said decomposition. In Sec. IV we describe the model used for type-II SPDC as well as the underlying assumptions used in its derivation. In Sec. IV.2, we go over numerical results for the Bloch-Messiah decomposition of certain waveguided squeezers (shown in Fig. 2) and show the difference (or lack thereof) between the input and output Schmidt modes to motivate the need to understand how these differences arise for different configurations. In Sec. V we develop a method to obtain analytic relations between input and output modes from the Bloch-Messiah decomposition of the discretized equations of motion by using symmetries of the underlying matrix which governs the evolution. We build our solution by first considering the SGVM regime for an unapodized squeezer in Sec. V.1. We then consider the simplest poling which gives rise to QPM in Sec. V.2 and move on to arbitrary poling (i.e. apodized poling to give rise to a Gaussian phase-matching function) in Sec. V.3. In Sec. V.4 we consider the double pass structure of Fig. 2(c). In Sec. V.5, we consider more general conditions relaxing the SGVM constraint. Table 3 concisely summarizes which assumptions are made for each section. Finally, in Sec. VI we conclude with our findings.
@@ III Bloch-Messiah decomposition
which represents a set of pairs of single-mode squeezing operations with squeezing parameters $\{r_{1},\ldots,r_{N}\}$ Cariolaro and Pierobon (2016a, b); McCutcheon (2018); Horoshko et al. (2019). As we are interested in two-mode squeezing, we need to convert the single-mode squeezing operations into two-mode squeezing operations to properly describe the output state. We can convert two single-mode squeezing operations into a two-mode squeezing operation, by combining them with a $50:50$ symmetric beamsplitter given by
@@ IV Model of twin-beam generation
We use the twin-beam dynamical equations derived in Ref.  Quesada et al., 2020 and used in Ref.  Houde and Quesada, 2023. We briefly described the underlying assumptions and express the relevant equations of motion. Without loss of generality, we assume that the beams propagate in the $z$ direction. We assume that each beam, $j=P,S,I$ for pump, signal, and idler respectively, has a linear dispersion around a central frequency $\bar{\omega}_{j}=\omega_{j,\bar{k}_{j}}$, which is associated with a central wavevector $\bar{k}_{j}$, such that
$\omega_{j,n}\approx\bar{\omega}_{j}+v_{j}(k-\bar{k}_{j}).$ (26)
@@ IV.1 Solving the equations of motion
which allows us to re-write the equations of motion (Eqs. 29, 30) in block-matrix form
@@
Figure 3: Temporal mode structure obtained from numerical Bloch-Messiah decomposition. Parameters for both structures were chosen to give the same number of signal photons $\langle N_{S}\rangle=10.5$. Plots are shown for poling function which gives rise to a Gaussian phase-matching function, however, the same phenomenological results are obtained for unpoled and periodic poling as well as for arbitrary levels of gain. a) First input Schmidt mode for single pass structures. b) First output Schmidt mode for single pass structures. We see that the input and output modes for the single pass structures have opposite phases and are therefore different. c) First input Schmidt mode for the double pass structure. d) First output Schmidt mode for double pass structure. We see that for the double pass structure the phases remain the same and so the input and output Schmidt modes are the same.
@@ IV.2 Numerical Bloch-Messiah results
Numerically, one can implement the Bloch-Messiah decomposition and study the input and output Schmidt modes for a given set of parameters. Reference Houde et al. (2024) goes through the numerical implementation in details which is included in the Python library TheWalrus Gupt, Izaac, and Quesada (2019a, b). We first begin by obtaining the solutions for the propagator using the Python library NeedALightHoude and Quesada (2022, 2023) and study its Bloch-Messiah decomposition.
@@
We then break the frequencies into a grid of $N=501$. For the double pass structure, the nonlinear regions are of the same length and the polarization swap is implemented by swapping the signal and idler velocities. For experimental purposes, the numerics can correspond to a pump with a central wavelength of 776 nm, and degenerate central frequencies of the signal and idler beams of 1552 nm, and a duration of  200 fs. We also tune the interactions strengths to generate $\langle N_{S}\rangle=\langle N_{I}\rangle=10.5$ photons in both configurations for signal and idler beams.
Figure 3 shows the different input and output modes for both single and double pass structures. Note that as outlined inHoude and Quesada (2023), we gauge away the free propagating phases and further choose a gauge where the imaginary part of the Schmidt modes is zero at the origin to produce these figures. In this gauge, we see that the input and output modes for the single pass configuration are not the same. Indeed, their phase structures are opposites. This difference is also seen for a single pass structure with no poling or QPM poling but has been left out for brevity. From this, we gather that a single pass structure does not give a perfect inline squeezer. The double pass structure, on the other hand, gives rise to input and output modes that are identical and is thus a perfect inline squeezer. Again these results hold for a double pass structure with no poling or QPM.
@@ V.1 Single pass without poling
where now $\bm{G}_{n,m}\leavevmode\nobreak\ =\leavevmode\nobreak\ \kappa\left(\omega_{n}-\bar{\omega}\right)\delta_{n,m}$ with $\bar{\omega}\equiv\bar{\omega}_{S}=\bar{\omega}_{I}$ and $\bm{F}_{n,m}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\gamma}{\sqrt{2\pi}}\beta_{P}(\omega_{n}+\omega_{m})\Delta\omega$. The solution to the propagator is $\bm{S}\leavevmode\nobreak\ =\leavevmode\nobreak\ \text{exp}\left(\Delta z\tilde{\bm{Q}}\right)$. In general, this matrix exponential cannot be evaluated analytically and as such we cannot obtain an analytical relations between the input and output modes from the Bloch-Messiah decomposition. However, with certain manipulations and using matrix properties, it is possible to obtain a form which readily factors into a Bloch-Messiah decomposition.
@@
The product $\bm{X}e^{\Delta z\bm{A}}$ is thus a real-symmetric matrix. Following the same arguments, we also find that $\bm{X}e^{-\Delta z\bm{A}^{T}}$ is also a real-symmetric matrix. Using the fact that $\bm{X}^{2}=\mathbb{1}_{N\times N}$, we can express the propagator as
$\bm{S}$ $=\bm{B}\begin{pmatrix}e^{\Delta z\bm{A}}&0_{2N\times 2N}\\ 0_{2N\times 2N}&e^{-\Delta z\bm{A}^{T}}\\ \end{pmatrix}\bm{B}^{T}$ (101)
@@ V.1.1 Frequency symmetric pump
$\bm{J}_{2N\times 2N}\bm{A}\bm{J}_{2N\times 2N}$ $=\begin{pmatrix}-\bm{J}_{N\times N}\bm{F}\bm{J}_{N\times N}&-\bm{J}_{N\times N}\bm{G}\bm{J}_{N\times N}\\ \bm{J}_{N\times N}\bm{G}\bm{J}_{N\times N}&-\bm{J}_{N\times N}\bm{F}\bm{J}_{N\times N}\\ \end{pmatrix}$ (165)
@@
By comparing the first vectors of each, we see that both the real and imaginary parts are flips of each other. We also see that for one vector, the real and imaginary parts are also flips of each other. This is not quite the behaviour shown in Fig. 3, however, recall that for plotting purposes we chose to be in a gauge where the imaginary part of the Schmidt mode is 00 at the origin. Furthermore, the free-propagation phases are also removed in these plots. Numerically, if one obtains the Schmidt modes above and applies the proper rotations to be in the same gauge, we obtain the results presented in Fig. 3.
@@ V.2 Single pass with quasi-phase-matching
where the -- subscript denotes that the poling function is negative. For the full propagator, we assume without loss of generality that the that the poling is symmetric (i.e. we pole with an odd number of domains) and we begin(and end) with positive poling. For a large number of domains, the solution should not depend on ±$\pm 1$ domains at the beginning or end of the waveguide. In this case, absorbing $\Delta z$ into $\bm{A}$, the propagator takes the form
@@
and therefore $\bm{X}e^{\bm{A}}e^{-\bm{A}^{T}}\cdots e^{-\bm{A}^{T}}e^{\bm{A}}$ is symmetric. All the previous arguments for the single poled case also apply here and so everything else follows for the Schmidt modes. We now wish to consider more arbitrary poling for more exotic phase-matching functions.

### v2

@@ I Introduction
Squeezed states of light are ubiquitous in quantum information and quantum computing in the optical domain Braunstein and van Loock (2005); Weedbrook et al. (2012); Braunstein (2005). Recently, there have been great advancements made on how to generate and use these states as resources in quantum computing both theoretically Christ et al. (2013, 2014); Helt and Quesada (2020); Quesada et al. (2022, 2020); Vernon et al. (2019); Arzani, Fabre, and Treps (2018); Houde and Quesada (2023); Grier et al. (2022) and experimentally Zhong et al. (2020, 2021, 2019); Eckstein et al. (2011); Wang et al. (2019); Deshpande et al. (2022); Arrazola et al. (2021); Triginer et al. (2020); Harder et al. (2013, 2016); Yanagimoto et al. (2023); Nehra et al. (2022); Thekkadath et al. (2024).
@@
One can gain information on both sets of modes by applying a Bloch-Messiah decomposition to the Heisenberg-picture propagator describing the evolution of the squeezer Cariolaro and Pierobon (2016a, b); McCutcheon (2018); Horoshko et al. (2019). This type of decomposition is readily implemented numerically Houde, McCutcheon, and Quesada (2024), however, it is generally very hard to obtain analytically, especially when a large number of modes are used to approximate the quasi-continuum encountered in waveguide geometries. In this paper, we look at different waveguided twin-beam squeezers, which are experimentally relevant Zhong et al. (2020, 2021, 2019), and develop a method for obtaining analytical relations between the input and output modes, without computing them explicitly. By working with discretized equations of motion and using the symmetries of the matrices describing the evolution, we are able to obtain analytical relations between the input and output modes from the Bloch-Messiah decomposition of the Heisenberg-picture propagator.
Figure 1: Schematic representation of a perfect inline squeezer. (a) Temporal mode structure with respect to central frequency $\bar{\omega}$, $f_{0}(\omega)$, that is matched to the perfect inline squeezer. (b) Wigner distribution of the injected quantum state, taken to be a single Fock state, for the matched mode. (c) Temporal mode structure with respect to central frequency $\bar{\omega}$, $f_{n}(\omega)$, of an unmatched mode. (d) Wigner distribution of the injected quantum state, taken to be a Cat state, for the unmatched mode. (e) Temporal mode structure of the output state for the matched input state. Note that the structure remains the same. (f) Wigner distribution of the output state for the matched case. The output is now is a Squeezed Fock state. (g) Temporal mode structure of the output state for the unmatched input state. Note that the structure remains the same. (h) Wigner distribution of the output state for the unmatched case. The Wigner distribution remains unchanged. Only the matched input generates a squeezed output.
@@
Figure 2: Waveguide poling configurations ($g(z)$) considered for the Bloch-Messiah decompositions of type-II SPDC. a) Unpoled single pass structure of length $z$, where the pump pulse enters from the left and the squeezed light exits to the right. b) Poled single pass structure where the nonlinear region is domain engineered(represented by up/down arrows). The domains can be chosen such that $z_{1}=z_{2}=\cdots=z_{n}$ to generate quasi-phase-matching or more arbitrary to generate more exotic phase-matching functions (e.g. Gaussian). c) Poled double pass structure where after the first pass, the beams are reflected back into the sample by reflecting off a mirror. The polarizations of the signal and idler are also swapped by passing through a quarter-wave plate (QWP) twice. Similar to the poled single pass, the values for the $z_{i}$’s can be chosen to generate quasi-phase-matching or more exotic phase-matching functions. In all configurations we assume no reflections at any of the interfaces and no propagation losses.
In this paper we focus on non-degenerate waveguided squeezers, however, one could also consider other architectures such as resonators and degenerate squeezers. In Table 1 we summarize known results on the spectral purity and mode-matching of degenerate ring-like and waveguided structures. It has been shown that ring-like structures can be spectrally pure, however, the input and output modes are not matchedQuesada et al. (2022), just as ring-like structures that have no nonlinear response. Therefore, ring-like structures are not good candidates for perfect inline squeezers. As for degenerate waveguided squeezers, it has been shown that they are highly spectrally impure when the linear dispersion dominatesHelt and Quesada (2020). It remains to be seen if such squeezers can be mode-matched, however, simply based on the spectral purity they are not good candidates for perfect inline squeezers. In Table 2 we summarize results related to this criteria but for non-degenerate ring-like and waveguided structures. The same results and conclusions hold for non-degenerate ring-like structures. As mentioned previously, non-degenerate waveguided sources can be spectrally pure and, as we show in this paper, can also be mode-matched therefore giving rise to a perfect inline squeezer. Finally, we note that if beamsplitter operations are easy to access between the two beams of the twin-beam (as is the case for frequency degenerate type-II, for which they correspond to polarizing beamsplitters) then it is easy to show that a perfect inline squeezer can be turned into two single-mode squeezing operations acting in parallel on the two modes Kalajdzievski and Quesada (2021).
@@
The paper is organized as follows. In Sec. II we provide a detailed analysis of why it is important to match input and output modes when squeezing states other than the vacuum. In Sec. III, we give a brief description of the Bloch-Messiah decomposition as well as how to identify the input and output Schmidt modes from said decomposition. In Sec. IV we describe the model used for type-II SPDC as well as the underlying assumptions used in its derivation. In Sec. IV.2, we go over numerical results for the Bloch-Messiah decomposition of certain waveguided squeezers (shown in Fig. 2) and show the difference (or lack thereof) between the input and output Schmidt modes to motivate the need to understand how these differences arise for different configurations. In Sec. V we develop a method to obtain analytic relations between input and output modes from the Bloch-Messiah decomposition of the discretized equations of motion by using symmetries of the underlying matrix which governs the evolution. We build our solution by first considering the SGVM regime for an unapodized squeezer in Sec. V.1. We then consider the simplest poling which gives rise to QPM in Sec. V.2 and move on to arbitrary poling (i.e. apodized poling to give rise to a Gaussian phase-matching function) in Sec. V.3. In Sec. V.4 we consider the double pass structure of Fig. 2(c). In Sec. V.4.1, we numerically consider how having distinct passes with different levels of gain affects the double pass result. In Sec. V.5, we consider more general conditions relaxing the SGVM constraint. Table 3 concisely summarizes which assumptions are made for each section. Finally, in Sec. VI we conclude with our findings.
@@ III Bloch-Messiah decomposition
which represents a set of pairs of single-mode squeezing operations with squeezing parameters $\{r_{1},\ldots,r_{N}\}$ Cariolaro and Pierobon (2016a, b); McCutcheon (2018); Horoshko et al. (2019). As we are interested in two-mode squeezing, we need to convert the single-mode squeezing operations into two-mode squeezing operations to properly describe the output state.
We can convert two single-mode squeezing operations into a two-mode squeezing operation, by combining them with a $50:50$ symmetric beamsplitter given by
@@ IV Model of twin-beam generation
We use the twin-beam dynamical equations derived in Ref.  Quesada et al., 2020 and used in Ref.  Houde and Quesada, 2023. We briefly describe the underlying assumptions and express the relevant equations of motion.
Without loss of generality, we assume that the beams propagate in the $z$ direction. We assume that each beam, $j=P,S,I$ for pump, signal, and idler respectively, has a linear dispersion around a central frequency $\bar{\omega}_{j}=\omega_{j,\bar{k}_{j}}$, which is associated with a central wavevector $\bar{k}_{j}$, such that
$\omega_{j,k}\approx\bar{\omega}_{j}+v_{j}(k-\bar{k}_{j}).$ (20)
@@ IV.1 Solving the equations of motion
which allows us to re-write the equations of motion (Eqs. 23, 24) in block-matrix form
@@
Figure 3: Temporal mode structure obtained from numerical Bloch-Messiah decomposition. Parameters for both structures were chosen to give the same average number of signal photons $\langle N_{S}\rangle=10.5$, giving rise to different Schmidt modes. Plots are shown for poling function which gives rise to a Gaussian phase-matching function, however, the same phenomenological results are obtained for unpoled and periodic poling as well as for arbitrary levels of gain. a) First input Schmidt mode for single pass structures. b) First output Schmidt mode for single pass structures. We see that the input and output modes for the single pass structures have opposite phases and are therefore different. c) First input Schmidt mode for the double pass structure. d) First output Schmidt mode for double pass structure. We see that for the double pass structure the phases remain the same and so the input and output Schmidt modes are the same.
@@ IV.2 Numerical Bloch-Messiah results
Numerically, one can implement the Bloch-Messiah decomposition and study the input and output Schmidt modes for a given set of parameters. Reference Houde, McCutcheon, and Quesada (2024) goes through the numerical implementation in details which is included in the Python library TheWalrus Gupt, Izaac, and Quesada (2019a, b). We first begin by obtaining the solutions for the propagator using the Python library NeedALightHoude and Quesada (2022, 2023) and study its Bloch-Messiah decomposition.
@@
We then break the frequencies into a grid of $N=501$. For the double pass structure, the nonlinear regions are of the same length and the polarization swap is implemented by swapping the signal and idler velocities. For experimental purposes, the numerics can correspond to a pump with a central wavelength of 776 nm, and degenerate central frequencies of the signal and idler beams of 1552 nm, and a duration of  200 fs. We also tune the interaction strengths to generate $\langle N_{S}\rangle=\langle N_{I}\rangle=10.5$ photons in both configurations for signal and idler beams.
Figure 3 shows the different input and output modes for both single and double pass structures. Note that as outlined inQuesada et al. (2020), we gauge away the free propagating phases and further choose a gauge where the imaginary part of the Schmidt modes is zero at the origin to produce these figures. In this gauge, we see that the input and output modes for the single pass configuration are not the same. Indeed, their phase structures are opposites. This difference is also seen for a single pass structure with no poling or QPM poling but has been left out for brevity. From this, we gather that a single pass structure does not give a perfect inline squeezer. The double pass structure, on the other hand, gives rise to input and output modes that are identical and is thus a perfect inline squeezer. Again these results hold for a double pass structure with no poling or QPM.
@@ V.1 Single pass without poling
where now $\bm{G}_{n,m}\leavevmode\nobreak\ =\leavevmode\nobreak\ \kappa\left(\omega_{n}-\bar{\omega}\right)\delta_{n,m}$ with $\bar{\omega}\equiv\bar{\omega}_{S}=\bar{\omega}_{I}$ and $\bm{F}_{n,m}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\gamma}{\sqrt{2\pi}}\beta_{P}(\omega_{n}+\omega_{m})\Delta\omega$. The solution to the propagator is $\bm{S}\leavevmode\nobreak\ =\leavevmode\nobreak\ \text{exp}\left(\Delta z\tilde{\bm{Q}}\right)$. In general, this matrix exponential cannot be evaluated analytically and as such we cannot obtain an analytical relation between the input and output modes from the Bloch-Messiah decomposition. However, with certain manipulations and using matrix properties, it is possible to obtain a form which readily factors into a Bloch-Messiah decomposition.
@@
The product $\bm{X}e^{\Delta z\bm{A}}$ is thus a real-symmetric matrix. Following the same arguments, we also find that $\bm{X}e^{-\Delta z\bm{A}^{T}}$ is also a real-symmetric matrix. Using the fact that $\bm{X}^{2}=\mathbb{1}_{2N\times 2N}$, we can express the propagator as
$\bm{S}$ $=\bm{B}\begin{pmatrix}e^{\Delta z\bm{A}}&0_{2N\times 2N}\\ 0_{2N\times 2N}&e^{-\Delta z\bm{A}^{T}}\\ \end{pmatrix}\bm{B}^{T}$
@@ V.1.1 Frequency symmetric pump
$\bm{J}_{2N\times 2N}\bm{A}\bm{J}_{2N\times 2N}$ $=\begin{pmatrix}-\bm{J}_{N\times N}\bm{F}\bm{J}_{N\times N}&-\bm{J}_{N\times N}\bm{G}\bm{J}_{N\times N}\\ \bm{J}_{N\times N}\bm{G}\bm{J}_{N\times N}&-\bm{J}_{N\times N}\bm{F}\bm{J}_{N\times N}\\ \end{pmatrix}$
@@
By comparing the first vectors of each, we see that both the real and imaginary parts are flips of each other. We also see that for one vector, the real and imaginary parts are also flips of each other. This is not quite the behaviour shown in Fig. 3, however, recall that for plotting purposes we chose to be in a gauge where the imaginary part of the Schmidt mode is 00 at the origin. Furthermore, the free-propagation phases are also removed in these plots(as outlined in reference Quesada et al. (2020)). Numerically, if one obtains the Schmidt modes above and applies the proper rotations to be in the same gauge, we obtain the results presented in Fig. 3.
@@ V.2 Single pass with quasi-phase-matching
where the -- subscript denotes that the poling function is negative. For the full propagator, we assume without loss of generality that the poling is symmetric (i.e. we pole with an odd number of domains) and we begin(and end) with positive poling. For a large number of domains, the solution should not depend on ±$\pm 1$ domains at the beginning or end of the waveguide. In this case, absorbing $\Delta z$ into $\bm{A}$, the propagator takes the form
@@
and therefore $\bm{X}e^{\bm{A}}e^{-\bm{A}^{T}}\cdots e^{-\bm{A}^{T}}e^{\bm{A}}$ is symmetric. All the previous arguments for the single unpoled case also apply here and so everything else follows for the Schmidt modes. We now wish to consider more arbitrary poling for more exotic phase-matching functions.